(c) Use the following information (in rupees):
Income (Y) = 1,00,000
Nominal Money Supply (M) = 80,000
Price Level (P) = 20
Calculate the money growth rate required to
finance the budget deficit of Rs.10,000 in an
economy. 4
1
Expert's answer
2021-07-26T17:09:02-0400

"Y=100,000"

"Money Supply( M)=80,000"

"Price Level (P)=20"

Money Growth Rate = "\frac {P\times Y}{M}"

"=\frac{20\times100,000}{80,000}"

"=25%" %